OpenStudy (brittanydosey):

Your parents are planning to rent you an apartment when you go to college. The annual contract is $635.00/month with a 2 month penalty for breaking the lease. The month-to-month contract is $715.00/month. What is the difference in the amount paid if you leave after 8 months?

OpenStudy (brittanydosey):

Multiply the monthly amount of the annual contract by the sum of the number of months you are there and the amount of the penalty. Put this number aside. Next, multiply the amount for the month-to-month contract by the number of months that you are there. Now subtract the number you put aside and the number that you just found.

OpenStudy (brittanydosey):

@kelliegirl33

OpenStudy (anonymous):

635(8) + 635(2) = 5080 + 1270 = 6350 715 (8) = 5720 6350 - 5720 = 630